# Segmentation-free Goodness of Pronunciation

### v2

@@ III Methods
To achieve these goals we propose two methods that will be detailed Section III-A and III-B.
@@ III-A Self-alignment GOP (GOP-SA)
Firstly, we consider the problem of mismatch between the segmentation of speech into phonetic units and the activations of the models used for GOP estimation as mentioned in Section II-B. This phenomenon is exemplified in Figure 2 and is especially critical for CTC trained models (right plots) which exhibit peaky activations for each phoneme that may not correspond in time with the corresponding speech segment. In order to address the problem we propose to use the same GOP definition as for DNN models (Eq. 3) but to perform the alignment of the target segment ($t_{1},t_{2}$) based on the same model used for GOP evaluation instead of an external forced aligner. We refer to this method as self-alignment GOP (GOP-SA). Figure 2 shows this alignment with red dashed vertical lines for the CTC-Loss trained models. We want to stress that the goal of alignment in this method is not to find the segment corresponding to the target phone, but, rather, to find the activations of the model corresponding to the target phone.
@@
III-B Alignment-free GOP (GOP-AF)
The second method is alignment-free (AF) and evaluates the GOP for a particular speech segment without the need for explicitly segmenting the utterance under test. The motivation for this method is to overcome the following limitations of standard GOP: 1) the evaluation of pronunciation of each phoneme is exclusively based on the corresponding speech segment (see Figure 1); 2) the evaluation of GOP is sensitive to the specific alignment ($t_{1},t_{2}$) and does not take into account uncertainty in alignment; 3) the common implementation of GOP uses Viterbi decoding and therefore only considers one path through the model possibly leaving out part of the probability mass; finally, 4) standard GOP does not allow for insertion and deletion erros.
We assume that we have recorded an utterance $O_{1}^{T}=\{o_{1},\dots,o_{T}\}$ with a canonical transcription $L_{\text{C}}=\{l_{1},\dots,l_{i},\dots,l_{|L_{\text{C}}|}\}$. We are interested in evaluating the pronunciation of a phoneme $l_{i}$. We can therefore split the canonical transcription into three parts, the left context ($L_{\text{L}}$) the target phoneme $l_{i}$, and the right context ($L_{\text{R}}$):
$L_{\text{C}}=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{l_{i},}_{\text{target phoneme}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}}\}.$
As in previous sections, we define $t_{1}$ and $t_{2}$ as the start and end frame index for the target phoneme $l_{i}$.
Figure 3: The figure exemplifies the computation of GOP-AF (Eq. 12) using a CTC-trained model. In the example, $\mathcal{L}_{\text{CTC}}(L_{\text{C}},O)$ can be computed using the graph at the top, whereas $\mathcal{L}_{\text{CTC}}(L_{\text{SDI}},O)$ using the graph at the bottom. For the sake of simplicity we omit the skip connections and self-loop connections.
Instead of committing to a specific alignment as in standard GOP, in the proposed alignment-free GOP (GOP-AF) we compute the log posterior for the target phoneme $l_{i}$ given the full observation sequence $O_{1}^{T}$ and all the phonemes in the left and right context:
[…]
This is the same definition that we introduced in [cao24b_interspeech], although not explicitly written as in Eq. 5, there.
@@ III-B Alignment-free GOP (GOP-AF)
[…]
@@
In the following subsections, we give theoretical and practical accounts on how to estimate the numerator and denominator in Eqs. 5 and 6.
III-C Alignment-free target phoneme posterior estimation
@@
We now consider a specific alignment $(t_{1},t_{2})$ for the target phoneme $l_{i}$ and we define the set of all possible alignments as
$\mathcal{A}(l_{i})=\{(t_{1},t_{2}):i\leq t_{1}\leq t_{2}\leq T-|L_{\text{C}}|+i\},$ (8)
where the lower bound for $t_{1}$ and the upper bound for $t_{2}$ ensure that there is at least one frame for each phoneme in the left context ($L_{\text{L}}$) and right context ($L_{\text{R}}$) of $l_{i}$.
@@
Special attention must be paid to the term $p(t_{1},t_{2}|O_{1}^{T})$ within both sums at the numerator and denominator of Eq. 9. This is the probability of a certain alignment for the $i$th segment, given the observation sequence, but independent of the actual transcription. We can interpret this as a prior with respect to the transcription over all possible segmentations. This distribution could be estimated from the training data. However, in our definition of GOP-AF, we make the simplifying assumption that this distribution is uniform. Under this assumption $p(t_{1},t_{2}|O_{1}^{T})$ is a constant and can be taken out of the sums and finally cancels out between the numerator and the denominator of Eq. 9 which becomes:
@@
Finally, the terms within the sums in Eq. 10 can be computed with the CTC’s forward variables $\alpha$ and $\beta$ as defined in [CTC]:
@@
In fact, the numerator in Eq. 11 is equivalent to the original definition of the CTC loss for the canonical pronunciation $L_{\text{C}}$ and the acoustic features $O_{1}^{T}$, except for a $-\log(.)$ term. Similarly, the denominator can be computed with CTC loss if we define a modified canonical pronunciation $L_{\text{SDI}}$ as:
$L_{\text{SDI}}$ =$=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{.*,}_{\text{any phoneme sequence}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}},\}.$
This is the set of all possible transcriptions in which the left and right contexts are equal to the canonical transcription, but we allow any sequence of phonemes in place of the target phoneme $l_{i}$ [^3] In this and later expressions, we have borrowed notation from regular expressions where “.” corresponds to any phoneme, “*” is zero or more occurrences, and “?” is zero or one occurrence.. This corresponds to allowing any number of “substitution”, “deletion” and “insertion” errors in pronunciation, justifying the SDI subscript. This CTC loss can be implemented by defining a graph exemplified in Figure 3 (bottom) and running the forward algorithm using this graph.
In summary, combining Eqs 5 and 11, the GOP-AF can be efficiently computed as the difference between the CTC loss for the pair ($L_{\text{SDI}}$,$O_{1}^{T}$) and the CTC loss for ($L_{\text{C}}$, $O_{1}^{T}$):
@@
Note that, because we are considering contributions from all possible segmentations, our method is able to deal with uncertainty in alignment. Also, the definition of $L_{\text{SDI}}$ with a graph as in Figure 3 gives us flexibility on the kind of pronunciation errors we consider. If the full graph is used, any substitution (S), deletion (D) and insertion (I) errors are considered. We refer to this version of the method as GOP-AF-SDI. If we remove the blue path from the graph, we only consider substitutions and deletions (GOP-AF-SD). This corresponds to a modified canonical pronunciation
$L_{\text{SD}}$ =$=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{.?,}_{\text{any phoneme or empty}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}},\}.$
If we also remove the red path, we only consider substitutions as in the traditional GOP (GOP-AF-S) and the modified canonical pronunciation is
$L_{S}$ =$=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{.,}_{\text{any phoneme}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}},\}.$
In [cao24b_interspeech] we used the unnormalized forward variables $\alpha$s to compute $\mathcal{L}_{CTC}(L_{\text{SDI}},O_{1}^{T})$ due of implementation issues. However, this may lead to numerical underflow errors when multiplying probabilities over longer speech segments. In this paper, we implement a version of the algorithm that uses the normalized forward variables $\hat{\alpha}_{t}(s)$ defined in [CTC] for both terms in Eq. 12. When making comparisons, we refer to this version as GOP-AF-numerical.
@@
III-D Alignment-free activation length estimation
@@
We call $\mathcal{N}_{\text{C}}$ the set of central nodes (yellow and purple) in the graph in Figure 3 that correspond to the “.*” term in $L_{\text{SDI}}$. Then, the expected value of the duration of the activations corresponding to $l_{i}$, is the sum over the whole observation sequence of the normalized forward variables $\hat{\alpha}$ corresponding to the nodes in $\mathcal{N}_{\text{C}}$:
@@
This expression can be computed efficiently with the forward algorithm that is also used to estimate the posterior of the target phoneme $l_{i}$ detailed in the previous section. Note that the definition of $L_{\text{SDI}}$ and $L_{\text{SD}}$ allow deletion of the target segment as a possible mispronunciation error. In this case, $\mathbb{E}[t_{2}-t_{1}|O_{1}^{T},L_{\text{L}},L_{\text{R}}]$ will tend to zero and GOP-AF will tend to diverge. In the actual implementation, we therefor floor this estimation by 1.
We expect this normalization factor to be most relevant for models that are not peaky in time. Peaky models tend to have activations that span over a single frame, thus making GOP-AF roughly equivalent to GOP-AF-Norm.
@@ III-E Computational complexity
We can compute the complexity of calculating GOP-AF (Eq. 12) with the help of dynamic programming. The graph used to compute $\mathcal{L}(L_{\text{C}},O_{1}^{T})$ (Figure 3, top) contains $|G|=2|L_{\text{C}}|+1$ nodes (the transcription labels interleaved by $\phi$s). The graph used to compute $\mathcal{L}(L_{\text{SDI}},O_{1}^{T})$ (Figure 3, bottom) contains $|G|=2|L_{\text{C}}|+|V|$ nodes, where $V$ is the set of output symbols from the neural network. The complexity of running dynamic programming on these graphs is therefore $O(T\times|G|)=O(T\times(|L_{\text{C}}|+|V|))$, that is, it is linear both in the length of the utterances, in the number of symbols in the canonical transcription augmented by the number of output symbols, which is usually a constant. For GOP-AF-Norm, the only overhead is the summation in Eq 13 because the normalized forward variables $\hat{\alpha}$ are already computed.
III-F Alignment-free GOP features
Similarly to what was done in [weigthed-GOP, improved_DNN, WEI2009896], the GOP-AF of the $i$th phoneme in a canonical sequence can be expanded into a feature vector using LPP (log posterior probability) and LPR (log posterior ratio vector):
@@ III-F Alignment-free GOP features
where $\text{LPP}=\log\mathit{p}(L_{\text{C}}|O_{1}^{T})=-\mathcal{L}_{\text{CTC}}(L_{\text{C}})$ follows the definition of CTC’s log-posterior in Eq.(4) and LPR is a vector of length $|C|+1$:
$\textbf{LPR}(i)=\left\{\log\left(\frac{\mathit{p}(L_{\text{C}}|O_{1}^{T}}{p(L|O_{1}^{T})}\right)\text{ with }L\in L_{\text{SD}}(l_{i})\right\}$ (15)
where $L_{\text{SD}}(l_{i})$ has already been defined in Section III-C. We do not include insertion errors here because that would result in infinite length for the feature vectors. Similar to GOP-AF-Norm, we append the expected activation count in Eq. 13 as an extra dimension of the feature vector which forms:
[…]
@@ III-G Measures of Peakiness
In Section II-B, we have introduced the peaky behavior of CTC-trained models both with respect to time (POT) and symbols (POS). Our two methods were introduced in part to cope with this behavior. Due to the definitions given in Section III, we expect the performance of GOP-SA (self-alignment) to increase with both POT and POS, in contrast to standard methods. We also expect GOP-AF (alignment-free) to be less affected by POT and POS. In order to test the effect of POS and POT on the different GOP methods in a reproducible way, we define objective measures of POS and POT.
@@
For POS, we borrow the idea from [CTC-ME] and define the average conditional entropy (ConEn) as:
@@ IV-A Data
CMU Kids [CMU_kids] comprises 9.1 hours of speech from children in the range from six to eleven years old, and total of 5180 sentences spoken by 65 females and 24 males. Child MDD is a more challenging task than adult MDD due to children’s heterogeneous characteristics in different language development stages [cao23_interspeech]. In CMU Kids, each utterance is equipped with a phonetic transcription which is firstly automatically generated and then corrected by humans if there is any mispronunciation (we refer to this as human transcription later in this paper). We obtain the canonical sequence for CMU Kids using the CMU Pronouncing dictionary provided by Kaldi [Povey_ASRU2011]. With the help of Levenshtein distance between the two transcriptions, we are able to label each phone in the canonical sequence with 1 (missing or mispronounced) or 0 (correctly pronounced). Note that this procedure is performed only for those utterances that are annotated as problematic. For other utterances, on the other hand, all the canonical phones are always labeled as 0. We refer to MDD experiments that use the labels obtained this way as “real errors”.
Following [cao23_interspeech], we also create an alternative MDD task based on the CMU Kids data, where we systematically change each phoneme in the canonical pronunciation to any other phoneme. We call this task “simulated errors” because we pretend that the recorded speech was incorrectly pronounced.
Finally, the Speechocean762 dataset [speechocean762] includes 5000 utterances read by 250 gender-balanced L2 English learners, half of which are children in the range from 5 years- to 15 years-old. The dataset is collected specifically for pronunciation assessment tasks with rich annotations at different linguistic levels. We focus on the phoneme-level where each phoneme is labeled as one of the three categories: 0 (mispronounced), 1 (strongly accented) and 2 (correctly spoken). Speechocean762 provides canonical phone-level transcriptions using the same phone inventory as the CMU pronunciation dictionary that we used for CMU Kids.
IV-B Baseline alignment model
The external segmentation used for all the experiments in this study is based on the same baseline alignment model. The model is a context independent GMM-HMM model that is trained using the standard Kaldi recipe “train-mono.sh” on the 100 hours training set “train-100-clean” from Librispeech [libri]. Since the output symbols of the baseline alignment model match the symbols of the canonical sequences, used the Kaldi’s command “gmm-align” for obtaining the segmentation for both CMU Kids and Speechocean762 for further experiments.

### v3

@@ III Methods
To achieve these goals we propose two methods that will be detailed in Section III-A and III-B.
@@ III-A Self-alignment GOP (GOP-SA)
Firstly, we consider the problem of mismatch between the segmentation of speech into phonetic units and the activations of the models used for GOP estimation as mentioned in Section II-B. This phenomenon is exemplified in Figure 2 and is especially critical for CTC trained models (right plots) which exhibit peaky activations for each phoneme that may not correspond in time with the corresponding speech segment. In order to address the problem, we propose to use the same GOP definition as for DNN models (Eq. 3) but to perform the alignment of the target segment ($t_{1},t_{2}$) based on the same model used for GOP evaluation instead of an external forced aligner. We refer to this method as self-alignment GOP (GOP-SA). Figure 2 shows this alignment with red dashed vertical lines for the CTC-Loss trained models. We want to stress that the goal of alignment in this method is not to find the segment corresponding to the target phone, but, rather, to find the activations of the model corresponding to the target phone.
@@
III-B Segmentation-free GOP (GOP-SF)
The second method is segmentation-free (SF) and evaluates the GOP for a particular speech segment without the need for explicitly segmenting the utterance under test. The motivation for this method is to overcome the following limitations of standard GOP: 1) the evaluation of pronunciation of each phoneme is exclusively based on the corresponding speech segment (see Figure 1); 2) the evaluation of GOP is sensitive to the specific alignment ($t_{1},t_{2}$) and does not take into account uncertainty in alignment; 3) the common implementation of GOP uses Viterbi decoding and therefore only considers one path through the model possibly leaving out part of the probability mass; finally, 4) standard GOP does not allow for insertion and deletion errors.
We assume that we have recorded an utterance $O_{1}^{T}=\{o_{1},\dots,o_{T}\}$ with a canonical transcription $L_{\text{C}}=\{l_{1},\dots,l_{i},\dots,l_{|L_{\text{C}}|}\}$. We are interested in evaluating the pronunciation of a phone $l_{i}$. We can therefore split the canonical transcription into three parts, the left context ($L_{\text{L}}$) the target phone $l_{i}$, and the right context ($L_{\text{R}}$):
$L_{\text{C}}=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{l_{i},}_{\text{target phone}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}}\}.$
As in previous sections, we define $t_{1}$ and $t_{2}$ as the start and end frame index for the target phone $l_{i}$.
Figure 3: The figure exemplifies the computation of GOP-SF (Eq. 12) using a CTC-trained model. In the example, $\mathcal{L}_{\text{CTC}}(L_{\text{C}},O_{1}^{T})$ can be computed using the graph at the top, whereas $\mathcal{L}_{\text{CTC}}(L_{\text{SDI}},O_{1}^{T})$ using the graph at the bottom. For the sake of simplicity we omit the skip connections and self-loop connections.
Instead of committing to a specific segmentation as in standard GOP, in the proposed segmentation-free GOP (GOP-SF) we compute the log posterior for the target phone $l_{i}$ given the full observation sequence $O_{1}^{T}$ and all the phonemes in the left and right context:
$\text{GOP-SF}(l_{i})=\log\left(p(l_{i}|O_{1}^{T},L_{\text{L}},L_{\text{R}})\right).$ (5)
This is the same definition that we introduced in [4], although not explicitly written as in Eq. 5, there.
@@ III-B Segmentation-free GOP (GOP-SF)
$\text{GOP-SF-Norm}(l_{i})=\frac{\log\left(p(l_{i}|O_{1}^{T},L_{\text{L}},L_{\text{R}})\right)}{\mathbb{E}[t_{2}-t_{1}|O_{1}^{T},L_{\text{L}},L_{\text{R}}]}.$ (6)
@@
In the following subsections, we give theoretical and practical accounts of how to estimate the numerator and denominator in Eqs. 5 and 6.
III-C Segmentation-free target phoneme posterior estimation
@@
We now consider a specific alignment $(t_{1},t_{2})$ for the target phone $l_{i}$ and we define the set of all possible alignments as
$\mathcal{A}(l_{i})=\{(t_{1},t_{2}):i-1<t_{1}\leq t_{2}<T-(|L_{\text{C}}|-i)\}.$ (8)
Because the left context $L_{\text{L}}$ and the right context $L_{\text{R}}$ contain respectively i1i-1 and $|L_{C}|-i$ symbols, the lower bound for $t_{1}$ and the upper bound for $t_{2}$ ensure that there is at least one frame for each phone in the left and right context.
@@
Special attention must be paid to the term $p(t_{1},t_{2}|O_{1}^{T})$ within both sums at the numerator and denominator of Eq. 9. This is the probability of a certain alignment for the $i$th segment, given the observation sequence, but independent of the actual transcription. We can interpret this as a prior with respect to the transcription over all possible segmentations. This distribution could be estimated from the training data. However, in our definition of GOP-SF, we make the simplifying assumption that this distribution is uniform. Under this assumption $p(t_{1},t_{2}|O_{1}^{T})$ is a constant and can be taken out of the sums and finally cancels out between the numerator and the denominator of Eq. 9 which becomes:
@@
Finally, the terms within the sums in Eq. 10 can be computed with the CTC's forward variables $\alpha$ and $\beta$ as defined in [14]:
@@
In fact, the numerator in Eq. 11 is equivalent to the original definition of the CTC loss for the canonical pronunciation $L_{\text{C}}$ and the acoustic features $O_{1}^{T}$, except for a $\log(.)$ term. Similarly, the denominator can be computed with CTC loss if we define a modified canonical pronunciation $L_{\text{SDI}}$ as:
$L_{\text{SDI}}$ $=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{.*,}_{\text{any phoneme sequence}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}},\}.$
This is the set of all possible transcriptions in which the left and right contexts are equal to the canonical transcription, but we allow any sequence of phonemes in place of the target phoneme $l_{i}$ [^4] In this and later expressions, we have borrowed notation from regular expressions where “.” corresponds to any phoneme, “*” is zero or more occurrences, and “?” is zero or one occurrence.. This corresponds to allowing any number of ``substitution'', ``deletion'' and ``insertion'' errors in pronunciation, justifying the SDI subscript. This CTC loss can be implemented by defining a graph exemplified in Figure 3 (bottom) and running the forward algorithm using this graph.
In summary, combining Eqs 5 and 11, the GOP-SF can be efficiently computed as the difference between the CTC loss for the pair ($L_{\text{SDI}}$$O_{1}^{T}$) and the CTC loss for ($L_{\text{C}}$, $O_{1}^{T}$):
$\text{GOP-SF}(l_{i})=\mathcal{L}_{CTC}(L_{\text{SDI}},O_{1}^{T})-\mathcal{L}_{CTC}(L_{\text{C}},O_{1}^{T}).$ (12)
Note that, because we are considering contributions from all possible segmentations, our method is able to deal with uncertainty in alignment. Also, the definition of $L_{\text{SDI}}$ with a graph as in Figure 3 gives us flexibility on the kind of pronunciation errors we consider. If the full graph is used, any substitution (S), deletion (D) and insertion (I) errors are considered. We refer to this version of the method as GOP-SF-SDI. If we remove the blue path from the graph, we only consider substitutions and deletions (GOP-SF-SD). This corresponds to a modified canonical pronunciation
$L_{\text{SD}}$ $=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{.?,}_{\text{any phoneme or empty}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}},\}.$
If we also remove the red path, we only consider substitutions as in the traditional GOP (GOP-SF-S) and the modified canonical pronunciation is
$L_{S}$ $=\{\underbrace{l_{1},\dots,l_{i-1}}_{L_{\text{L}}},\underbrace{.,}_{\text{any phoneme}}\underbrace{l_{i+1},\dots,l_{|L_{\text{C}}|}}_{L_{\text{R}}},\}.$
@@
III-D Segmentation-free activation length estimation
@@
We call $\mathcal{N}_{\text{C}}$ the set of central nodes (yellow and purple) in the graph in Figure 3 that correspond to the ``.*'' term in $L_{\text{SDI}}$. Then, the expected value of the duration of the activations corresponding to $l_{i}$, is the sum over the whole observation sequence of the normalized forward variables $\hat{\alpha}$ corresponding to the nodes in $\mathcal{N}_{\text{C}}$:
@@
This expression can be efficiently computed with the forward algorithm that is also used to estimate the posterior of the target phoneme $l_{i}$ detailed in the previous section. Note that the definitions of $L_{\text{SDI}}$ and $L_{\text{SD}}$ allow deletion of the target segment as a possible mispronunciation error. In this case, $\mathbb{E}[t_{2}-t_{1}|O_{1}^{T},L_{\text{L}},L_{\text{R}}]$ will tend to zero and GOP-SF will tend to diverge. In the actual implementation, therefore, we define a floor value of 1.
We expect this normalization factor to be most relevant for models that are not peaky in time. Peaky models tend to have activations that span over a single frame, thus making GOP-SF roughly equivalent to GOP-SF-Norm.
@@ III-E Computational complexity
III-F Segmentation-free GOP features
Similarly to the approaches in [7, 20, 36], the GOP-SF of the $i$th phoneme in a canonical sequence can be expanded into a feature vector using LPP (log posterior probability) and LPR (log posterior ratio vector):
$\textbf{FGOP-SF}(l_{i})=\{\text{LPP},\textbf{LPR}(l_{i})\}$ (14)
where $\text{LPP}=\log\mathit{p}(L_{\text{C}}|O_{1}^{T})=-\mathcal{L}_{\text{CTC}}(L_{\text{C}})$ follows the definition of CTC's log-posterior in Eq.(4) and LPR is a vector of length $|C|+1$:
$\textbf{LPR}(i)=\left\{\log\left(\frac{\mathit{p}(L_{\text{C}}|O_{1}^{T})}{p(L|O_{1}^{T})}\right)\text{ with }L\in L_{\text{SD}}(l_{i})\right\}$ (15)
where $L_{\text{SD}}(l_{i})$ has already been defined in Section III-C. We do not include insertion errors here because that would result in infinite length for the feature vectors. Similar to GOP-SF-Norm, we append the expected activation count in Eq. 13 as an extra dimension of the feature vector which forms:
$\textbf{FGOP-SF-Norm}(l_{i})=\{\text{LPP},\textbf{LPR}(l_{i}),\text{Occ}(i)\}.$ (16)
@@ III-G Measures of Peakiness
In Section II-B, we have introduced the peaky behavior of CTC-trained models both with respect to time (POT) and symbols (POS). Our two methods were introduced in part to cope with this behavior. Due to the definitions given in Section III, we expect the performance of GOP-SA (self-alignment) to increase with both POT and POS, in contrast to standard methods. We also expect GOP-SF (segmentation-free) to be less affected by POT and POS. In order to test the effect of POS and POT on the different GOP methods in a reproducible way, we define objective measures of POS and POT.
@@
For POS, we borrow the idea from [26] and define the average conditional entropy (ConEn) as:
@@ IV-A Data
CMU Kids [9] comprises 9.1 hours of speech from children in the range 6–11 years old, and a total of 5180 sentences spoken by 65 females and 24 males. In CMU Kids, each utterance is equipped with a phonetic transcription and utterances including mispronunciations are marked. We determine the distribution of pronunciation errors by comparing [^5] The comparing procedure is performed only for those utterances that are marked as problematic. For the correct utterances, on the other hand, we simply label all the canonical phonemes as correct. the phonetic transcription with the canonical pronunciation from the CMU pronunciation dictionary provided by Kaldi [30]. There are 150899 labeled phonemes in total, among which, 90.2% are correct; 3.2% are substitution errors; 2.3% are deletion errors; and 4.3% are insertion errors. We refer to MDD experiments that use the labels obtained this way as ``real errors''.
Following [3], we also create an alternative MDD task based on the CMU Kids data, where we systematically change each phoneme in the canonical pronunciation to any other phoneme. We call this task ``simulated errors'' because we pretend that the recorded speech was incorrectly pronounced.
Finally, the speechocean762 dataset [40] includes 5000 utterances read by 250 gender-balanced L2 English learners, half of which are children in the range from 5 years- to 15 years-old. The dataset is collected specifically for pronunciation assessment tasks with rich annotations at different linguistic levels. We focus on the phoneme-level where each phoneme is labeled as one of the three categories: 0 (mispronounced), 1 (strongly accented) and 2 (correctly spoken). speechocean762 provides canonical phone-level transcriptions using the same phone inventory as the CMU pronunciation dictionary that we used for CMU Kids. We were also interested in comparing the results for children and adults. In order to do this we split the speechocean762 test set according to the reported age of the speaker. The children (age 6–11) account for 38.4% of the utterances whereas adults (age 12–43) account for 61.6%. Note however, that the results on speechocean762 children are not comparable to those in CMU Kids because the two data sets define different tasks (ternary vs binary assessment).
IV-B Baseline segmentation model
The external segmentation used for all the experiments in this study is based on the same baseline segmentation model. The model is a context-independent GMM-HMM model that is trained using the standard Kaldi recipe “train-mono.sh” on the 100 hours training set “train-100-clean” from Librispeech [29]. Since the output symbols of the baseline segmentation model match the symbols of the canonical sequences, we used the Kaldi's command ``gmm-align'' for obtaining the segmentation for both CMU Kids and speechocean762 for further experiments.